OpenStudy (anonymous):

What is the sum 7/2+(-3/5)

OpenStudy (hba):

Take lcm.

OpenStudy (aravindg):

well take lcm :)

OpenStudy (aravindg):

did u get the answer @Lila.hasipi ?

OpenStudy (anonymous):

lcm??

OpenStudy (aravindg):

yep :)

OpenStudy (hba):

@Lila.hasipi Please do not try to swap for answers on a particular test.

OpenStudy (anonymous):

Idk what lcm means

OpenStudy (aravindg):

\[\large \frac{a}{b} + \frac{c}{d} = \frac{ad+bc}{bd}\]

OpenStudy (aravindg):

use this !!

OpenStudy (hba):

You multiply the denominator that is your lcm in this ques. See @AravindG has shown.

OpenStudy (anonymous):

k

OpenStudy (hba):

So start working.

OpenStudy (anonymous):

LOL SURE :))))